Jabu had a few marbles. Today he played and double his number of marbles. Then Thabo gave him three marbles for free. Jabu has 21 marbles now. How many did he have before he started playing?
1
Expert's answer
2020-07-06T19:04:16-0400

Let "x" denote the initial number of marbles.

Today Jabu played and double his number of marbles


"2x"

Then Thabo gave him three marbles for free


"2x+3"

Jabu has 21 marbles now


"2x+3=21"

Solve for "x"


"2x+3-3=21-3"

"2x=18"

"{2x\over 2}={18\over 2}"

"x=9"

Jabu had 9 marbles before he started playing.